# Non-Adaptive Multi-Stage Algorithm and Bounds for Group Testing with Prior Statistics

### v4

@@ Abstract
In this paper, we propose an efficient multi-stage algorithm for non-adaptive Group Testing (GT) with general correlated prior statistics. The proposed solution can be applied to any correlated statistical prior represented in trellis, e.g., finite state machines and Markov processes. We introduce a variation of List Viterbi Algorithm (LVA) to enable accurate recovery using much fewer tests than objectives, which efficiently gains from the correlated prior statistics structure. We also provide a sufficiency bound to the number of pooled tests required by any Maximum A Posteriori (MAP) decoder with an arbitrary correlation between infected items. Our numerical results demonstrate that the proposed Multi-Stage GT (MSGT) algorithm can obtain the optimal MAP performance with feasible complexity in practical regimes, such as with COVID-19 and sparse signal recovery applications, and reduce in the scenarios tested the number of pooled tests by at least $25\%$ compared to existing classical low complexity GT algorithms. Moreover, we analytically characterize the complexity of the proposed MSGT algorithm that guarantees its efficiency.
@@ I Introduction
Classical Group Testing (GT) aims to detect a small number of “defective” items within a large population by mixing samples into as few pooled tests as possible. The idea of GT was first introduced during World War II when it was necessary to discover soldiers infected with Syphilis. Dorfman [1] showed that the required number of tests could be reduced if multiple blood samples were grouped into pools. When the samples that participate in the next pool are selected iteratively based on the previous pool test results, the GT algorithm is called adaptive. In contrast, in non-adaptive GT, the whole process is designed in advance. Since it was first suggested, the GT problem has been investigated and generalized to many areas and applications, among them disease detection [2, 3, 4], cyber security applications [5], pattern matching algorithms [6] and communication [7, 8].
@@
All of these applications imply a strong connection between GT and Compressed Sensing (CS) as two methods for sparse signal recovery that share common applications [9, 10, 11, 12]. The main difference between the two is that CS aims to recover a real-valued signal [13], while GT recovers a binary signal [14] or discrete-values [15, 16, 4]. Thus, one can consider GT as a Boolean CS [17, 18].
@@
Traditional GT and its performance (i.e., the tradeoff between the number of pool tests and recovery algorithm complexity), focuses on the probabilistic model in which the items are identically distributed and independent [19, 14]. Recent research explores cases where prior information about the correlation of objects is available [20, 2, 3, 21, 22]. The motivation for this approach arises from the fact that correlated prior statistics have the potential to achieve higher recovery rates and reduce the number of required tests. In disease detection, leveraging information about the proximity between individuals, represented by contact tracing information or graphs, can lead to significant savings in pool tests [20, 2, 3, 21]. However, previous GT works presented solutions designed for specific models and applications and may not be extended easily to other models and applications. In numerous signal processing applications, correlation between different frequencies, time signals, or among different sensors can also be utilized to achieve more precise estimations [23, 13, 24]. Hidden Markov Model (HMM) is a common model for many physical signals, such as speech signals [25], human motion [26], and spectrum occupancy in communication systems [27]. Infections can be also modeled as HMMs [28]. For example, [29] presents a GT solution for a specific HMM derived from contact tracing. To the best of our knowledge, no existing solution addresses the GT problem with general Markovian priors and applicable to a wide range of diverse applications.
@@
In this work, we introduce a practical non-adaptive Multi-Stage Group Testing (MSGT) algorithm for correlated items with prior statistics. The proposed multi-stage algorithm employs a new variation of the parallel List Viterbi Algorithm (LVA) [30, 31, 32] we designed for GT to enable accurate low complexity recovery using fewer tests. The proposed algorithm can be applied for any statistical prior represented in trellis [33], e.g., finite-state machines and Markov processes. Using LVA, MSGT leverages those statistics to estimate the defective set efficiently, even in a regime below the maximum likelihood (ML) upper bound. Furthermore, we show how the algorithm’s parameters can be tuned to achieve a maximum probability of success without exceeding the limitation of the available computational capacity. We derive a lower bound that considers the exact priors of the problem and provides analytical results that characterize the MSGT computational complexity efficiency. We provide an analytical sufficiency bound for the number of pooled tests needed by any maximum a-posteriori (MAP) decoder. This bound holds for any correlation between infected items. It also applies to multiple access communication systems, where $K$ out of $N$ users transmit information simultaneously. This is considered in [34, 35, 15] without prior statistical information. In contrast, existing analyses with prior statistics assume all $N$ users transmit at the same time [36, 37, 38, 39, 40, 41, 42]. Our numerical results demonstrate that in practical regimes for COVID-19 [43, 44] and sparse signal recovery in signal processing [9, 10, 11, 13], the low computational complexity MSGT algorithm proposed herein can reduce in the scenarios tested the number of pool tests by at least $25\%$.
@@
The rest of this paper is organized as follows. Section II formally describes the GT model with correlated prior statistics. Section III presents the MSGT algorithm and the analytical results, and Section IV presents a sufficiency bound for MAP decoder. Section V details the simulation evaluation. Finally, Section VI provides concluding remarks and future directions.
@@ II Problem Formulation
Given a set of individuals $\mathcal{N}$, the objective in GT is to discover a small subset $\mathcal{K}$ of unknown defective items using the minimum number of measurements $T$. Let $N=|\mathcal{N}|$, $K=|\mathcal{K}|$ denote the total number of items and the number of defective items, respectively, where $K=\mathcal{O}(N^{\theta})$ for some $\theta\in\left[0,1\right)$. The binary vector $\mathbf{U}\in\left\{0,1\right\}^{N}$ represents the population, such that $U_{i}=1$ indicates that the $i$-th item is defective. We assume that the set of the individuals is sparse, such that $\theta\leq 1/3$[17, 19, 45, 14], and that each item can be in one of 2 states: defective and not-defective. For each item, there is some prior probability that it is defective, $\left\{\mathbf{\pi}_{i}\right\}_{i=1}^{N}\in\left[0,1\right]^{N}$, $\pi_{i}=P\left(U_{i}=1\right)$. The correlation between the state of the $i$-th item and the states of the previous $\tau$ items is represented by $\mathbf{\Phi}_{i}\in\left[0,1\right]^{2^{\tau}\times 2}$. $\tau=1,2,\dots$ represents the number of memory steps of the underlying process considered in the recovery stage. When there is no memory in the process, $\tau=0$ and $\{\pi_{i}\}_{i=1}^{N}$ hold all the prior information. $\mathbf{\Phi}_{i}[l,k]=P\left(U_{i}=s_{k}\lvert\left(U_{i-\tau},...,U_{i-1}\right)=s_{l}\right)$, where $s_{l}$ is the binary representation of $l$ by a $\tau$-length binary vector and $s_{k}\in\{0,1\}$ (for example, see Fig. 7(b)).
@@
$Y_{i}=\bigvee_{j\in\mathcal{K}}X_{i}(j)=\bigvee_{j=1}^{N}X_{i}(j)U_{j}$
where $\bigvee$ is the Boolean OR function. All vectors are column vectors and the $\log$ operator denotes $\log_{2}$, unless stated otherwise. These definitions are illustrated in Fig. 1.
@@
Given $\mathbf{X}$ and the outcome vector $Y$, the recovery success criterion in GT can be measured using various metrics [14]. The main metrics we will use herein are exact recovery and partial recovery. In terms of exact recovery, the goal is to detect the precise subset of defective items $\mathcal{K}$. Therefore, given the estimated defective set $\hat{\mathcal{K}}=\hat{\mathcal{K}}(N,K,\mathbf{X},Y)$, we define the average error probability by [^1] For simplicity of notation, $P_{s}$ and $P_{e}=1-P_{s}$ refer to success and error probabilities in the exact recovery analysis.
$P_{e}^{exact}\triangleq\frac{1}{\binom{N}{K}}\sum_{\mathcal{K}:\left|\mathcal{K}\right|=K}P\left(\hat{\mathcal{K}}\neq\mathcal{K}\right).$
In partial recovery, we allow a false positive (i.e., $|\hat{\mathcal{K}}\setminus\mathcal{K}|$) and false negative (i.e., $|\mathcal{K}\setminus\hat{\mathcal{K}}|$) detection rate. Thus, the average partial success rate is given by:
@@
$P_{s}^{partial}\triangleq\frac{1}{\binom{N}{K}}\sum_{\mathcal{K}:\left|\mathcal{K}\right|\neq K}\frac{\left|\hat{\mathcal{K}}\cap\mathcal{K}\right|}{K}.$
@@
In conclusion, the goal of non-adaptive GT with correlated prior data is to identify the $K$ defective items out of the $N$ items, where there is correlation between different infected items. To achieve this, one needs to design a $T\times N$ testing matrix and an efficient and practical recovery algorithm that can exploit correlated priors. The algorithm can identify the subset of infected items by observing $Y^{T}$ with high probability and with feasible computational complexity. Given the knowledge of the correlated prior data items and the available computational resources, one can design the testing matrix and a recovery algorithm to maximize the success probability.
@@ III Main Results
In this section, we introduce the efficient multi-stage recovery algorithm for any statistical prior represented in a trellis diagram [33], detailed in Algorithm 1. In the first stage, standard low-complexity algorithms [19] reduce the search space independently of prior correlations. This reduction is guaranteed by new analytical results we derive. In the second stage, the algorithm employs a novel adaptation of the LVA [32], designed for GT to enable accurate low-complexity recovery using fewer tests by exploiting the correlated prior information. Additionally, we derive a bound to ensure the low complexity of the entire algorithm. Section III-A describes the proposed algorithm. Section III-B provides analytical results, followed by a discussion in Section  III-C. For a detailed explanation of all the algorithms used as integral components of Algorithm 1, see Appendix A.

### v5

# arXiv:2402.10018v5[cs.IT] 28 Feb 2026

@@ Abstract
In this paper, we propose an efficient two-stage decoding algorithm for non-adaptive Group Testing (GT) with general correlated prior statistics. The proposed solution can be applied to any correlated statistical prior represented in trellis, e.g., finite state machines and Markov processes. We introduce a variation of List Viterbi Algorithm (LVA) to enable accurate recovery using much fewer tests than objectives, which efficiently gains from the correlated prior statistics structure. We also provide a sufficiency bound to the number of pooled tests required by any Maximum A Posteriori (MAP) decoder with an arbitrary correlation, i.e., dependence between infected items. Our numerical results demonstrate that the proposed two-stage decoding GT (2SDGT) algorithm can obtain the optimal MAP performance with feasible complexity in practical regimes, such as with COVID-19 and sparse signal recovery applications, and reduce in the scenarios tested the number of pooled tests by at least $25\%$ compared to existing classical low complexity GT algorithms. Moreover, we analytically characterize the complexity of the proposed 2SDGT algorithm that guarantees its efficiency.
@@ I Introduction
Classical Group Testing (GT) aims to detect a small number of “defective” items within a large population by mixing samples into as few pooled tests as possible. The idea of GT was first introduced during World War II when it was necessary to discover soldiers infected with Syphilis. Dorfman [25] showed that the required number of tests could be reduced if multiple blood samples were grouped into pools. When the samples that participate in the next pool are selected iteratively based on the previous pool test results, the GT algorithm is called adaptive. In contrast, in non-adaptive GT, the whole process is designed in advance. Since it was first suggested, the GT problem has been investigated and generalized to many areas and applications, among them disease detection [34, 61, 22, 60], cyber security applications [64], pattern matching algorithms [35] and communication [65, 20].
@@
All of these applications imply a strong connection between GT and Compressed Sensing as two methods for sparse signal recovery that share common applications [32, 63, 1, 47]. The main difference between the two is that Compressed Sensing aims to recover a real-valued signal [26], while GT recovers a binary signal [4] or discrete-values [18, 21, 22]. Thus, one can consider GT as a Boolean Compressed Sensing [8, 55].
@@
Traditional GT and its performance (i.e., the tradeoff between the number of pooled tests and recovery algorithm complexity), focuses on the probabilistic model in which the items are identically distributed and independent [3, 4]. Recent research explores cases where prior information about the correlation (i.e., dependence) of objects is available [48, 34, 61, 39, 6]. The motivation for this approach arises from the fact that correlated prior statistics have the potential to achieve higher recovery rates and reduce the number of required tests. In disease detection, leveraging information about the proximity between individuals, represented by contact tracing information or graphs, can lead to significant savings in pooled tests [48, 34, 61, 39]. However, previous GT works presented solutions designed for specific models and applications and may not extend easily to other models and applications. In numerous signal processing applications, correlation between different frequencies, time signals, or among different sensors can also be utilized to achieve more precise estimations [68, 26, 7]. Hidden Markov Model is a common model for many physical signals, such as speech signals [28], human motion [62], and spectrum occupancy in communication systems [31]. Infections can be also modeled as Hidden Markov Models [14]. For example, [13] presents a GT solution for a specific Hidden Markov Model derived from contact tracing. To the best of our knowledge, no existing solution addresses the GT problem with general Markovian priors and applicable to a wide range of diverse applications.
@@
In this work, we addresses the GT problem with general prior statistics. We adopt a more general formulation in which the number of defective items is not assumed to be fixed a priori, but is instead induced by the underlying correlated prior distribution. In particular, the number of defectives is treated as a random variable, and the system design is based on a typical cardinality interval of the number of defective items. We introduce a practical non-adaptive Two-Stage Decoding GT (2SDGT) algorithm for correlated items with prior statistics. The proposed two-stage decoding algorithm employs a new variation of the parallel List Viterbi Algorithm (LVA) [5, 43, 56] we designed for GT to enable accurate low complexity recovery using fewer tests. The proposed algorithm can be applied to general Markovian correlated priors, which may be represented using trellis structures [46], e.g., finite-state machines and Markov processes. While classical group testing assumes exchangeable items, many practical settings admit structured dependencies arising from spatial, temporal, or graph-based relationships. Rather than prescribing a specific physical model, this work adopts a general trellis-based prior as a unifying abstraction for correlated item statistics. This formulation subsumes the i.i.d. model as a special case and enables a single decoding framework to operate across diverse applications where correlation information is available or can be estimated. The use of the LVA is motivated by the observation that, under the Markovian prior model, the population sequence admits a natural trellis representation, and that approximate MAP decoding over this trellis can be performed efficiently by enumerating a small list of the most likely trajectories rather than a single path. Using LVA, 2SDGT leverages those statistics to estimate the defective set efficiently, even in a regime below the maximum likelihood (ML) upper bound. Furthermore, we show how the algorithm’s parameters can be tuned to achieve a maximum probability of success without exceeding the limitation of the available computational capacity. We derive a lower bound that considers the exact priors of the problem and provides analytical results that characterize the 2SDGT computational complexity efficiency. We provide an analytical sufficiency bound for the number of pooled tests needed by any maximum a-posteriori (MAP) decoder. This bound holds for any dependence between infected items. It also applies to multiple access communication systems, where $K$ out of $N$ users transmit information simultaneously. This is considered in [67, 66, 18] without prior statistical information. In contrast, existing analyses with prior statistics assume all $N$ users transmit at the same time [58, 69, 70, 12, 11, 51, 52]. Our numerical results demonstrate that in practical regimes for COVID-19 [44, 9] and sparse signal recovery in signal processing [32, 63, 1, 26], the low computational complexity 2SDGT algorithm proposed herein can reduce in the scenarios tested the number of pooled tests by at least $25\%$.
@@
The rest of this paper is organized as follows. Section II formally describes the GT model with correlated prior statistics. Section III presents the 2SDGT algorithm and the analytical results, and Section IV presents a sufficiency bound for MAP decoder. Section V details the simulation evaluation. Finally, Section VI provides concluding remarks and future directions.
@@ II Problem Formulation
Given a set of individuals $\mathcal{N}$, the objective in GT is to discover a small random subset $\mathcal{K}$ of unknown defective items using the minimum number of measurements $T$. Let $N=|\mathcal{N}|$ denote the total number of items. Let $\mathcal{U}=\{0,1\}$ be the state alphabet, and let $\mathbf{U}=(U_{1},\ldots,U_{N})\in\mathcal{U}^{N}$ represent the random population vector, where $U_{i}=1$ indicates that the $i$-th item is defective. Under this model, the random defective set is $\mathcal{K}\triangleq\{i\in\mathcal{N}:U_{i}=1\}$ and the number of defective items is the random variable $K\triangleq|\mathcal{K}|=\sum_{i=1}^{N}U_{i}$. We denote by $\mu\triangleq\mathbb{E}[K]=\sum_{i=1}^{N}\Pr(U_{i}=1)$ the expected number of defective items.
We model the prior over the sequence as a $\tau$-order Markov process. The joint law of the first $\tau$ variables is specified by an initial distribution $\pi(u_{1},\ldots,u_{\tau})\triangleq P(U_{1}=u_{1},\ldots,U_{\tau}=u_{\tau})$, for $(u_{1},\ldots,u_{\tau})\in\mathcal{U}^{\tau}$. This initial distribution generalizes the marginal parameters $\{\pi_{i}\}$ stated earlier: for $\tau=1$ it reduces to $\pi(u_{1})=P(U_{1}=u_{1})$ (equivalently, $\pi_{1}(u_{1})$), whereas for $\tau>1$ the model requires specifying the joint initialization of the first $\tau$ states. For $i\geq\tau+1$, the dependence on the previous $\tau$ states is specified by the transition kernels $\{\Phi_{i}\}_{i=\tau+1}^{N}$, defined by
$\Phi_{i}\!\left[u_{i}\mid u_{i-\tau},\ldots,u_{i-1}\right]\triangleq P\!\left(U_{i}=u_{i}\,\big|\,U_{i-1}=u_{i-1},\ldots,U_{i-\tau}=u_{i-\tau}\right).$
Equivalently, the transition of the $i$-th item can be represented in matrix form by
$\boldsymbol{\Phi}_{i}\in[0,1]^{2^{\tau}\times 2},\quad\text{with transition probability}\quad\boldsymbol{\Phi}_{i}[\ell,k]=P\!\big(U_{i}=s_{k}\,\big|\,(U_{i-\tau},\ldots,U_{i-1})=s_{\ell}\big),$
where $s_{\ell}\in\{0,1\}^{\tau}$ is the $\tau$-bit binary representation of $\ell\in\{0,\ldots,2^{\tau}-1\}$ and $s_{k}\in\{0,1\}$ (see Fig. 1). The distribution of the entire chain is then
$P\left(U_{1}=u_{1},\ldots,U_{N}=u_{N}\right)=\pi(u_{1},\ldots,u_{\tau})\prod_{i=\tau+1}^{N}\Phi_{i}\!\left[u_{i}\mid u_{i-\tau},\ldots,u_{i-1}\right].$
When there is no memory ($\tau=0$), the process reduces to independent (but not necessarily identical) marginals, and the prior information is fully captured by the marginal probabilities
$P(U_{1}=u_{1},\ldots,U_{N}=u_{N})=\prod_{i=1}^{N}P(U_{i}=u_{i}).$
Figure 1: Correlation (i.e., dependence of objects) matrix $\mathbf{\Phi}_{i}$ for $\tau=2$.
Instead of conditioning on a fixed realization of $K$ [4], we introduce design sparsity bounds $K_{\min}$ and $K_{\max}$ that are used for test-matrix design and algorithmic thresholds, and are not assumed to equal the realized value of $K$. A convenient choice is $K_{\min}\triangleq\lfloor(1-\varepsilon)\mu\rfloor$ and $K_{\max}\triangleq\lceil(1+\varepsilon)\mu\rceil$, for some $\varepsilon\in(0,1)$. In subsequent sections, the recovery performance is analyzed by conditioning on the typical event $\color[rgb]{0,0,0}\{K_{\min}\leq K\leq K_{\max}\}$}, while accounting separately for the tail probability $\Pr(K<K_{\min})+\Pr(K>K_{\max})$ via concentration bounds on $K$. However, it is important to emphasize that the random variable $K$ is not restricted a priori and may take any value in $\{0,\ldots,N\}$. We focus on sparse regimes with $\mu=\mathcal{O}(N^{\theta})$ for some $\theta\in[0,1)$, and in particular $\theta\leq 1/3$ [8, 3, 54, 4].
@@
$Y_{i}=\bigvee_{j\in\mathcal{K}}X_{i}(j)=\bigvee_{j=1}^{N}X_{i}(j)U_{j},$
where $\bigvee$ is the Boolean OR function.
Given $\mathbf{X}$ and the outcome vector $Y$, the recovery success criterion in GT can be measured using various metrics [4]. The main metrics we will use herein are exact recovery and partial recovery. We note that the trellis (Markov) prior adopted herein should be interpreted as a modeling abstraction rather than as an assumption of a fixed physical ordering of items. The index $i$ does not necessarily correspond to spatial or temporal adjacency; instead, it represents an ordering induced by latent structure such as clustering, spatial proximity, temporal evolution, graph embeddings, or other side information available to the designer. When the items are exchangeable and no correlation information is available, the proposed model naturally reduces to the independent prior as a special case ($\tau=0$). When correlations are present, an appropriate permutation of item indices allows the dependency structure to be represented using a low-order Markov model. Under such correlated priors, the test matrix need not be permutation-invariant, as exploiting prior information inherently breaks symmetry across items. This behavior is not a limitation of the method but rather a consequence of leveraging structured side information to improve recovery performance. The proposed framework does not assume that such priors are always observable in practice; instead, it provides a general recovery architecture capable of exploiting correlated priors whenever they are available or can be inferred from data.
We consider the prior probability of a defective set $\mathcal{K}$ through the population vector $\mathbf{U}=\mathbf{1}_{\mathcal{K}}$: $P(\mathcal{K})\ \triangleq\ P(\mathbf{U}=\mathbf{1}_{\mathcal{K}})$, where $P(\mathbf{U})$ is induced by the $\tau$-order Markov model described above in this section. Here $P(\mathcal{K})$ denotes the probability of the defective set induced by the prior distribution on the population vector $\mathbf{U}$, without conditioning on a fixed sparsity level. Note that this probability is defined over all subsets $\mathcal{K}\subseteq\mathcal{N}$ and does not restrict the cardinality $|\mathcal{K}|$.
@@
In terms of exact recovery, the goal is to detect the precise subset of defective items $\mathcal{K}$. Therefore, given the estimated defective set $\hat{\mathcal{K}}=\hat{\mathcal{K}}\!\left(N,K_{\min},K_{\max},\mathbf{X},Y,\mathbf{\pi},\{\mathbf{\Phi}_{i}\}_{i=\tau+1}^{N}\right)$, we define the average exact error probability by [^1] For simplicity of notation, $P_{s}$ and $P_{e}=1-P_{s}$ refer to success and error probabilities in the exact recovery analysis.
$P_{e}^{\mathrm{exact}}\triangleq\sum_{\mathcal{K}\subseteq\mathcal{N}}P(\mathcal{K})\;\Pr\!\left(\hat{\mathcal{K}}\neq\mathcal{K}\,\big|\,\mathcal{K}\right)=\Pr\!\left(\hat{\mathcal{K}}\neq\mathcal{K}\right).$
Moreover, for any design sparsity bounds $(K_{\min},K_{\max})$,
$P_{e}^{\mathrm{exact}}\leq\Pr\!\left(\hat{\mathcal{K}}\neq\mathcal{K}\mid{\color[rgb]{0,0,0}K_{\min}\leq K\leq K_{\max}}\right){\color[rgb]{0,0,0}+\Pr(K<K_{\min})+\Pr(K>K_{\max}).}$ (1)
Here $(K_{\min},K_{\max})$ are design parameters rather than the realized sparsity $K$.
[…]
@@
$P_{s}^{\mathrm{partial}}\triangleq\sum_{\mathcal{K}\subseteq\mathcal{N}}P(\mathcal{K})\;\mathbb{E}\left[\left.{|\hat{\mathcal{K}}\cap\mathcal{K}|}/{|\mathcal{K}|}\,\right|\mathcal{K}\right].$
@@
To conclude, for a sparse subset of infected items with random size $K$ (with $\mathbb{E}[K]=\mu$) out of $N$, the goal in non-adaptive GT with correlated prior data items is to design a $T\times N$ testing matrix and an efficient and practical recovery algorithm that can exploit correlated priors, such that by observing $Y^{T}$ we can identify the subset of infected items with high probability and with feasible computational complexity. Thus, given the knowledge of the correlated prior data items and the available computational resources, the test designer could design the testing matrix and a recovery algorithm to maximize the success probability.
@@ III Main Results
In this section, we introduce the efficient two-stage [^2] Although we refer to the decoding procedure as consisting of two stages, the testing process itself is fully non-adaptive: the considered testing matrix is fixed in advance (see Section III-A), and no additional tests are designed based on intermediate decoding results. The term “stage” is used here only to distinguish between successive computational steps of the proposed decoder herein. recovery algorithm for any statistical prior represented in a trellis diagram [46], detailed in Algorithm 1. In the first stage, standard low-complexity algorithms [3] reduce the search space independently of prior correlations. This reduction is guaranteed by new analytical results we derive. In the second stage, the algorithm employs a novel adaptation of the LVA [56], designed for GT to enable accurate low-complexity recovery using fewer tests by exploiting the correlated prior information. Additionally, we derive a bound to ensure the low complexity of the entire algorithm. Under the correlated prior (See Section II), the number of defectives $K=\sum_{i=1}^{N}U_{i}$ is a random variable. The algorithm is not assumed to know the realized $K$. Whenever a step requires a cardinality (filtering the LVA output and enumerating candidate supports), we use $(K_{\min},K_{\max})$ as the design bounds. Performance under random $K$ is handled by conditioning on the typical event $\{K_{\min}\leq K\leq K_{\max}\}$ and adding the tail probability $\Pr(K<K_{\min})+\Pr(K>K_{\max})$, as defined in (1) and elaborated in Section III-C. A key design parameter of the proposed 2SDGT algorithm presented in this section is the constant $\gamma\geq 1$, which controls the maximal candidate support size admitted from the LVA stage to the final MAP decoding stage. Specifically, only trajectories whose estimated number of defective items lies within the interval $K_{\min},\lfloor\gamma K_{\max}\rfloor]$ are retained for candidate generation. The parameter $\gamma$ therefore directly governs the tradeoff between computational complexity and recovery performance: larger values of $\gamma$ increase robustness to model mismatch and estimation errors at the cost of increased candidate enumeration, while smaller values enforce stricter complexity constraints. Section III-A describes the proposed algorithm. Section III-B provides analytical results, followed by a discussion in Section  III-C. For a detailed explanation of all the algorithms used as integral components of Algorithm 1, see Appendix A.